A particle with mass 4.0 kg is initially at rest at the origin (x=0). It then moves along the x-axis acted on by a force F(x)=18.0-2.0x2, where F is in Newtons if x is expressed in meters. Find the kinetic energy of the particle when it reaches x=3 m.

A Honda Civic (mass=1090 kg) is initially traveling due East at 18.3 m/s when it suddenly collides head on with a Toyota Camry (mass=1612 kg) initially traveling due West at 31.8 m/s. The collision is completely inelastic. Find the speed of the Civic after the collision.

An object moves along the x-axis, with it’s position as a function of time given by x(t)=0.7t3-6.2t, where the position x is given in meters if t is given in seconds. What is the speed of the particle at time t=4.36 s?
Answer #1

• Work done = change in krt- kEf-k.G; = 3 f dx K.EE -0 = k Ef = S3 (18 – 2 x ²)dn = (180- 2,333 K. Ef = 18x3 - 2x33 a 0 = 363kindly give rating

Post each question separately